two planes leave newark airport at the same time. The rate of speed of the northbound plane is 100 mils per hour.,than that of the southbound plane. at the end of 5 hours, the planes are 2000 miles apart. Find the rate of each?
Answer by mananth(16949)


let speed of south bound plane be x
north bound plane = x+100 ( presume it is 100 more)
Time thay travel = 5 hours
distance of south boundplane = 5x
distance of north boundplane = 5(x+100)
..
5x+5(x+100)=2000
5x+5x+500=2000
10x+500=2000
10x=2000-500
10x= 1500
x= 150 mph south bound plane
..
north bound plane speed = x+100 = 150+100=250 mph
--------------
CHECK
150*5+250*5= 2000